# Mitigating Outlier Activations in Low-Precision Fine-Tuning of Language Models

### 15 Dec 2023 v2

@@ 2.2 Handling Outliers in Low-precision Back-propagation
Over the past few years, low-precision training of deep learning models has gained popularity in reducing the training cost. For instance, FP16 mixed precision training [Micikevicius et al., 2017] is nowadays a commonplace methodology to fine-tune language models. Integer data type has also been extensively studied for back-propagation computations by [Zhang et al., 2020, Zhao et al., 2021, Zhu et al., 2020, Ghaffari et al., 2022]. Moreover, using higher-bit integer formats such as INT12 for both back-propagation and forward propagation is proposed by [Hosseini et al., 2023].
@@ 4.1 Information Loss in Low-precision Number Formats
$\mathrm{sens}_{\theta}(X)=\frac{(\mathrm{lev}_{\theta}(X))^{2}}{\mathbb{V}(X)}.$ (4)
@@ 4.2 Analysing Outlier Activations as a Mixture Distribution
INT8 Untreated Outliers 80.5 86.4 80.9 91.8 58.5 69.9/81.9 47.0
Table 1: Metric performance of integer fine-tuning of BERT on selected GLUE tasks. The reported metric for QQP and MRPC is accuracy and F1 score, for QNLI, MNLI, RTE, and SST-2 is accuracy, for STSB is the Pearson-Spearman correlation, and for CoLA is the Matthews correlation.

### 13 Jan 2024 v3

@@ 2.2 Handling Outliers in Low-precision Back-propagation
Over the past few years, low-precision training of deep learning models has gained popularity in reducing the training cost. For instance, FP16 mixed precision training [Micikevicius et al., 2017] is nowadays a commonplace methodology to fine-tune language models. Integer data type has also been extensively studied for back-propagation computations by [Zhang et al., 2020, Zhao et al., 2021, Zhu et al., 2020, Ghaffari et al., 2022]. Moreover, using higher-bit integer formats such as INT12 for both back-propagation and forward propagation is proposed by [Tayaranian et al., 2023].
@@ 4.1 Information Loss in Low-precision Number Formats
$\mathrm{sens}_{\theta}(X)=\frac{(\mathrm{lev}_{\theta}(X))^{2}}{\mathbb{V}(X)},$ (4)
where $\theta$ is the parameters of $X$ distribution.
@@ 4.2 Analysing Outlier Activations as a Mixture Distribution
Table 1: Metric performance of integer fine-tuning of BERT on selected GLUE tasks. The reported metric for MRPC is accuracy and F1 score, for QNLI, MNLI, RTE, and SST-2 is accuracy, for STSB is the Pearson-Spearman correlation, and for CoLA is the Matthews correlation.
@@
INT8 Untreated Outliers 80.9 86.4 80.9 91.8 58.5 69.9/81.9 43.5